OpenStudy (anonymous):

Please help :( 15^-4 ------ 15^5

OpenStudy (anonymous):

@SheldonEinstein

OpenStudy (anonymous):

@gaby6 @surdawi

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@bigboss123

OpenStudy (anonymous):

Just subtract the powers.. :)

OpenStudy (anonymous):

@Miyuru I know but wouldnt it be 15^-9 its not a answer Help please

OpenStudy (anonymous):

\[\frac{ a^b }{ a^c }=a^{b-c}\]

OpenStudy (anonymous):

15^-9 is correct

OpenStudy (anonymous):

Well what I get is this...\[\frac{15^{-4}}{15^5}=\frac{1}{15^{9}}\]

OpenStudy (anonymous):

15^-9 = 2.6*10^-11

OpenStudy (anonymous):

Which expression is equivalent to (7^3)-2? 1. 1 ------------- 7*7*7*7*7*7 2. 7 3. 1 ------------------- 6*6*6*6*6*6*6 4. 1 __ 7

OpenStudy (anonymous):

Well what do you get for the answer ?

OpenStudy (anonymous):

@Miyuru 1.